write the first five terms of the sequence
Tn = -2.22n + 1.3

Answer by Edwin McCravy(20059)

write the first five terms of the sequence
Tn = -2.22n + 1.3
write the first five terms of the sequence


This means to replace n by 1, then by 2, then by 3, then by 4, then by 5.

Tn = 2n+3

Replacing n by 1:

T1 = -2.22(1) + 1.3 = -2.22+1.3 = -.92

Replacing n by 2:

T2 = -2.22(2) + 1.3 = -4.44+1.3 = -3.14

Replacing n by 3:

T3 = -2.22(3) + 1.3 = -6.66+1.3 = -5.36

Replacing n by 4:

T4 = -2.22(4) + 1.3 = -8.88+1.3 = -7.58

Replacing n by 5:

T5 = -2.22(5) + 1.3 = -11.1+1.3 = -9.8

and the sequence is

T1, T2, T3, T4, T5

or  -.92, -3.14, -5.36, -7.58, -9.8

Edwin